OpenStudy (anonymous):

Find the partial derivatives of the function. w = sqrt(r^2+s^2+t^2) w/r = ? w/s = ? w/t = ?

OpenStudy (anonymous):

Use chain rule.

OpenStudy (anonymous):

Consider it like: \[ w=\sqrt u, \quad u=r^2+s^2+t^2 \]Now: \[ \frac{\partial w}{\partial r}=\frac{d w}{d u}\frac{\partial u}{\partial r} \]

OpenStudy (anonymous):

We can start with \[ w'(u) = (\sqrt u)'=\frac 1{2\sqrt u}=\frac 1{2\sqrt {r^2+s^2+t^2}} \]

OpenStudy (anonymous):

@planetmonkey Can you find the rest?

OpenStudy (anonymous):

You gotta find \[ \frac{\partial u}{\partial r},\frac{\partial u}{\partial s},\frac{\partial u}{\partial t} \]

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

can you just plz type up the full explanation?